OpenStudy (anonymous):

can someone explain how to graph: y=1/2*cos(3ø+m/3)

OpenStudy (xapproachesinfinity):

think of the parent function \(\ y=\cos \theta \)

OpenStudy (xapproachesinfinity):

what is the period of cos ?

OpenStudy (xapproachesinfinity):

The period is pi

OpenStudy (xapproachesinfinity):

2pi i mean sorry

OpenStudy (xapproachesinfinity):

so knowing that we want to find the period of \(y=\frac{1}{2} \cos (3\theta+\frac{m}{3}) \)

OpenStudy (xapproachesinfinity):

we set \(3\theta +\frac{m}{3}=2\pi\) solve for theta

OpenStudy (xapproachesinfinity):

you get \(\theta =\frac{2\pi}{3}-\frac{m}{9}\) this our period

OpenStudy (xapproachesinfinity):

so the period is where the graph repeats it self

OpenStudy (xapproachesinfinity):

what you need to think of is the graph of cos theta how you can translate it to that graph

OpenStudy (xapproachesinfinity):

1/2 is the amplitude period I did it horizontal shift solve theta /3 +m/3=0 this is good enough to tell you about the graph

OpenStudy (anonymous):

thank you so much, made so much sense